Two identical conducting spheres, fixed in place, attract each other with an electrostatic force of 0.121 N when their center-to-center separation is 68.6 cm. The spheres are then connected by a thin conducting wire. When the wire is removed, the spheres repel each other with an electrostatic force of 0.0358 N. Of the initial charges on the spheres, with a positive net charge, what was (a) the negative charge on one of them and (b) the positive charge on the other? (Assume the negative charge has smaller magnitude.)
Answer #1

Initial electrostatic force,

F = kq1q2 / r^2

0.121 = 9*109*q1*q2 / (0.686)2

q1q2 = 6.33*10-12

Since conducting spheres attract each, that means q1 and q2 have opposite charges.

so, q1q2 = - 6.33*10-12 ..........(1)

electrostatic force when spheres repel each other,

F = kQ^2 / r^2

When spheres are connected with a wire and when wire is removes, charge will be distributed evenly among all the spheres.

so, two spheres have equal charge of Q.

0.0358 = 9*109*Q^2 / (0.686)2

Q^2 = 1.87*10-12

Q = \pm 1.368*10-6 C

q1 + q2 = 2Q = \pm 2.737*10-6

q2 = 2.737*10-6 - q1 ........(2)

From eq (1) and (2),

q1 * ( 2.736*10-6 - q1) = - 6.33*10-12

q12 - 2.736*10-6*q1 - 6.33*10-12 = 0

By solving above quadratic equation,

q1 = -1.495*10-6 C , q2 = 4.23*10-6 C

Or q1 = 4.23*10-6 C , q2 =  -1.495*10-6 C

(a)

Negative charge on one of them = -1.495*10-6 C

(b)

Positive charge on the other = 4.23*10-6 C